Vous êtes sur la page 1sur 5

ETA BALANCE WORKS STATUS

MOM : ETA Balance Work dated 13-Mar-14 Attendee : ANC Mr Madhav, Mr.Siji Sr. No. Activities All Types of CP gratings ANC had instructed ETA to deliver all the CP gratings by 15-Jan-14. If not delivered same will be procured by ANC at ETA 's risk and cost. ETA stated that they are procuring the material and delivery status can be updated on 20Feb-14 after they pay the supplier. CP grating at basement Catch basin grating (600X600) FAHU room grating(300X300) Man hole cover (30x30) Pipe end cap near PHE pump room Action By Attendee : ETA Mr.Alok,Mr.Rahul ETA REPLY DATED 7TH MARCH'2014 ETA COMMITMENT DATES ON 7-Mar-14 START DATE Attendee : HAIL Mr.Anand END DATE ANC REVIEW ON 13-MAR-14

ETA

a b c d e

Tenant Ducting & Balance LCP's ETA Reply is same. Balance LCP's ANC had instructed ETA to deliver all the balance LCP's will be delivered by 08-March-14 and start the ducting works from 28th floor to and installation with cabling and downwards for the balance Tenant AHU rooms by 15-Jan- ETA and ANC termination shall be completed by 14. ETA stated that they will deliver the balance 64 30th of March'14. Ducting work is LCP'sand the delivery status will be updated on 20-Febbeing carried out by ANC. 14. It is found that there is obstruction by PPR line on the 13th Floor. Same will be rectified by ETA by 13th March'2014. On the 9th floor, there is obstruction by existing ducting to tenant AHU ducting. ANC is requested to modify the existing ducting thru their ducting sub-contractor already working for tenant AHUs. Materials will be delivered by 11th March-2014

10Nos 6Nos 3Nos 6Nos 2Nos

On the same day, when Materials are on site. Necessary necessary civil works on civil works to be carried out by basement floors to install gratings ANC as per Email sent to be carried out by ANC as per Email sent.

On the same day, when necessary civil works on basement floors to install gratings to be carried out by ANC as per Email sent.

ETA requested ANC to install the gratings supplied by them. ANC stated same will be confirmed after discussing with Mr.Mukesh

8th March'2014

30th March'2014

20 Panels delivered on 12-Mar-14 Balance 44 panels will be delivered on 20-Mar-14 & 25-Mar14 (22/lot) resepectively

ANC asked ETA to confirm their programme for Fire fighting lines rectification within the tenant AHU rooms. ETA replied that these works can be taken up only upon completion of CFO inspection and the start date will be tentatively 21-Feb-14.

ETA

PPR rectification on 13h Floor, start by 12th March'2014.

PPR rectification finish by 13th March'2014.

Color coding of PHE pipes. Installation of Fire signages & signages for busducts

ANC stated that existing duct installed by ETA hence ETA to rectify. But ETA request ANC to rectify the same due to lack of their ducting subcontractor. ANC stated same will be confirmed after discussing with Mr.Mukesh 19-Mar-14

ETA

12th March'2014

15th March'2014

1 OF 5

ETA BALANCE WORKS STATUS

Sr. No. Activities 5 Domestic & flushing tank piping (For receiving domestic & Flushing water from utility corridors. ETA confirmed that works have commenced today and shall be complete by 18-Feb-14. Electrical room shaft closure with Fire sealant ANC asked ETA to start the electrical room shaft closure with fire sealant. ETA stated that same will be done on priority and they have called M/s. HILTI for a meeting tomorrow. However, the works shall tentatively commence by 20-Feb-14. ANC asked ETA to do the earthing for ELV equipments. ETA stated that the same is not within their scope. ANC insisted that ETA has to do the same at the earliest and any issues related to scope shall be dealt as per subcontract conditions.

Action By ETA

ETA REPLY DATED 7TH MARCH'2014

START DATE

END DATE

ANC REVIEW ON 13-MAR-14

Done

NA

NA Done

ETA

NA

NA

NA Done

ANC

The ETA reply is same. ELV system is being carried by HAIL under a PO from ANC. ETA has completed all the earthing as per approved shop drawings.

NA

NA

ANC stated all 220V AC equipment earthings to be provided by ETA for ELV equipment. ETA stated that ELV earthling is not in their scope.

Fire pump exhaust connection re-routing

ETA

9 10 11

Fire pumps exhaust piping cladding Split AC installation at BMS room O & M Manuals re-submission after incorporating comments. ETA stated that it is in process and will be done by 25-Feb-14. ANC raised concerns on the delay in re-submission and instructed ETA to resubmit on priority.

ETA ETA ETA

Work is in progress. Rerouting will be completed by 8th March'2014. Insulation/cladding will be completed by 15th March'2014

13th March'2014

15th March'2014

Rerouting done.And End date is 19-Mar-14 Done NA NA

In -progress

1st March'2014

15th March'2014

Electrical submitted HVAC-15-Mar-14 PHE & FPS-19-Mar-14

12

CEI certification. ETA will update the status on 13-Feb-14

ETA

Permanent power drawing/plan approval is received & certification is already forwarded. High rise building drawing/plan is received & already forwarded.High rise certification is balance.

In-progress

10th March'2014

Completed

2 OF 5

ETA BALANCE WORKS STATUS

Sr. No. Activities 13 14 15 16 FIRE NOC certification. ETA to submit the required compliance certificated as required by CFO for final certificates. ANC asked ETA to submit the GF Service lobby FCU relocation drawing as per actual site conditions. ETA to submit modification drawing by 14-Feb- 14. Sump pump rectification at PHE pump room (1 No) Issues related to BMS Integration ANC asked ETA to resolve all the issues which are effecting integrations. ETA stated that same will be resolved

Action By ETA ETA ETA

ETA REPLY DATED 7TH MARCH'2014 Meeting required with GIFT CFO for High rise NOC certification. Rest is done. Done Rectification at vendors workshop in progress.

START DATE To be determined after meeting with GIFT CFO NA 13th March'2014

END DATE To be determined affer meeting with GIFT CFO NA 15th March'2014

ANC REVIEW ON 13-MAR-14 ETA to submit compliance certificates on 14-Mar-14

Done Final date will be confimed on 14-Mar-14

Fire damper logic programming rectification

ETA

The ETA reply is same.Work completed uptill 15th floor. For balance floors,work is in progress in co-ordination with Honeywell.

As per progress of HAIL

As per progress of HAIL In progress

Termination details of LT Panels & HT panels to be given to HAIL Diesel engine pump termination details to be given to HAIL

ETA

ETA

The ETA reply is same.ANC to schedule a meeting with Honeywell as ETA has fulfilled all the requirements given by Honeywell.

To be determined after meeting with Honeywell

To be determined after meeting with Honeywell

Already explained the requiement of HAIL to ETA..ETA stated that same will be done on 19Mar-14

Danfoss Engineer to be called for VFD integrations of chilled water pumps and tenant AHUs.

ETA

The ETA reply is same.Completed for chilled water To be done by Honeywell pumps. For tenant AHUs work progressively as per installation of shall be completed progressively LCP works start by 10th March'2014. with the installation of LCPs by 30th March'2014. Done The ETA reply is same.ANC to schedule a meeting with Honeywell as ETA has fulfilled all NA

To be done by Honeywell progressively as per installation of LCP works finish by 30th March'2014. DONE NA Done Already explained the requiement of HAIL to

e f

15th floor network switch power supply to be provided. Provision of potential free contact for sump pump - 4 Nos

ETA ETA

To be determined after meeting

To be determined after

3 OF 5

ETA BALANCE WORKS STATUS

Sr. No. Activities g Principle vendor to be called for BTU meter modbus details for tenant AHU integration

Action By ETA

ETA REPLY DATED 7TH MARCH'2014 Honeywell as ETA has fulfilled all the requirements given by Honeywell. The ETA reply is same.This matter was discussed with ANC/ PMC and ITUAL in a meeting on 14-02-14 and explained to all lighting controls are provided by ETA based on the scheme agreed thru shop drawings. HAIL was then asked by ITUAL to take up the matter appropriately.

START DATE To be determined after meeting with Honeywell

END DATE To be determined after meeting with Honeywell

ANC REVIEW ON 13-MAR-14 requiement of HAIL to ETA..ETA stated that same will be done on 19Mar-14

Electrical lighting control panel integration issue to be resolved

ANC/ HAIL

NA

NA

Noted 19-Mar-14 ETA asked 3 days access to all floors as they are charging the entire system. ANC stated same will be confirmed after discussing with Mr.Mukesh In progress 19-Mar-14 ETA stated that same will be installed by 22Mar-14 M/s.ETA stated that issue will be resolved on 20-Mar-14

Booster pump parameters

ETA

Flow switch commissioing through sprinkler

k l j k

FCU CFM cheking in all the three speeds Display of Incorrect readings of Energy meters during "OFF" condition 400 Dia BTU Meter installation at ETS room M/s.ANC stated that lift lobyy pressurization is being failed as required readings as per standands are not shown to PMC and asked M/s.ETA to resolve the issue Ms/ANC aksed M/s.ETA to instruct respective services engineers to be presnt during integration.M/s.ETA agreed for the same

Agreed by ETA

4 OF 5

ETA BALANCE WORKS STATUS

Sr. No. Activities 17 Balance as built drawings submission. ANC raised concern on delay and stated that this will delay handing over formalities. ANC asked ETA to provide the necessary clearance as agreed earlier for installation of window cleaning equipment at terrace. ETA stated that same will be provided after examining the site condition. ANC asked ETA to rectify the pre check/snag points given by client for basement. ETA stated that same will be attended.

Action By ETA

ETA REPLY DATED 7TH MARCH'2014 In -progress

START DATE 1st March'2014

END DATE 15th March'2014

ANC REVIEW ON 13-MAR-14

20-Mar-14 ANC has been requested to arrange for a meeting with window cleaning vendor.

18

ETA

19

ETA

Work in progress.

Separate sheet enclosed Noted

5 OF 5

Centres d'intérêt liés